# OmniACT: A Dataset and Benchmark for Enabling Multimodal Generalist Autonomous Agents for Desktop and Web

### v1

@@ Abstract
footnotetext: $\dagger$ These authors contributed equally. The order is determined by dice rolling.
For decades, human-computer interaction has fundamentally been manual. Even today, almost all productive work done on the computer necessitates human input at every step. Autonomous virtual agents represent an exciting step in automating many of these menial tasks. Virtual agents would empower users with limited technical proficiency to harness the full possibilities of computer systems. They could also enable the efficient streamlining of numerous computer tasks, ranging from calendar management to complex travel bookings, with minimal human intervention. In this paper, we introduce OmniACT, the first-of-a-kind dataset and benchmark for assessing an agent’s capability to generate executable programs to accomplish computer tasks. Our scope extends beyond traditional web automation, covering a diverse range of desktop applications. The dataset consists of fundamental tasks such as “Play the next song”, as well as longer horizon tasks such as “Send an email to John Doe mentioning the time and place to meet”. Specifically, given a pair of screen image and a visually-grounded natural language task, the goal is to generate a script capable of fully executing the task. We run several strong baseline language model agents on our benchmark. The strongest baseline, GPT-4, performs the best on our benchmark However, its performance level still reaches only 15% of the human proficiency in generating executable scripts capable of completing the task, demonstrating the challenge of our task for conventional web agents. Our benchmark provides a platform to measure and evaluate the progress of language model agents in automating computer tasks and motivates future work towards building multimodal models that bridge large language models and the visual grounding of computer screens.
Figure 1: OmniACT dataset and benchmark for enabling autonomous human-computer interaction agents. The left shows an image paired with a natural language task description as the input, the right shows the resulting action script to be executed on the screen. Examples are presented from Stocks, Apartments.com, and Weather application.
@@ 1 Introduction
Performing computer tasks based on natural language instructions has been a long-standing goal of artificial intelligence [45]. One concrete objective in the line of research is to develop generalist agents that can assist humans in doing computer tasks [19], such as “Order a pizza from Domino’s” or “Write a message to John.” The agent should be able to open the application and perform the task. Executing these actions on a personal computer involves a sequence of interactions with a mouse and keyboard. For example, the simple task of writing an email involves hovering over the application icon, clicking it, clicking the ‘New Email’ button, writing the content of the email, and clicking send. Successfully sending an email requires accurately predicting the correct action at each step and accurately executing it, which is a herculean task even for the best agents today [13].
@@
We evaluate several language model-based agent baselines on this dataset, including LLaMA [43], Vicuna [7], Palmyra-X (43B) [2], InstructPalmyra-30B [41], GPT 3.5, and GPT-4 [30]. We experiment with fine-tuning Vicuna-13B and LLaMA-13B models using QLoRA [10]. We also benchmark multimodal baseline LLaVa-v1.5-7B, LLaVa-v1.5-13B [43] and GPT-4-vision-preview [50] for the task. Our findings highlight the necessity for a multimodal model capable of executing these tasks, and our analysis provides insights into promising future work in the space. Our key contributions are outlined as follows:
@@ 2 Related Work
Table 1: Comparison of OmniACT with other related benchmarks.
@@ 2.1 UI Understanding
User interface (UI) understanding has garnered interest from researchers in the machine learning and human-computer interaction communities, evolving with various models focusing on understanding the semantics of mobile and web user interfaces. UIBert [3], PixelBERT [15], ActionBert [14], VUT [23], Screen2Words [44], WidgetCaptioning [22] and Pix2Act [36] are notable models in this area. They propose approaches for learning the user-interface semantics of the mobile screen using the image and view hierarchy. These models have demonstrated effectiveness in tasks like capability prediction, screen segmentation and understanding, and screen caption generation. Lexi [4] and Spotlight [20] propose models that use vision-only inputs to minimize the reliance on metadata such as view hierarchy. Furata et al. [11] demonstrates the use of fine-tuning for multimodal web navigation. The majority of machine learning models trained for UI understanding leverage the Rico dataset [8] and its extensions, which contain 64,462 unique Android screens and metadata. In addition, Banerjee et al. [4] released the UICaptions dataset, which consists of diverse image-captions pairs across a wide range of applications. PixelHelp [21] also released a corpus to train models that can interpret natural language instructions and map them to mobile user interface actions.
@@ 2.2 Autonomous Computer Agents
The advent of large language models (LLMs) has been pivotal in the rapid advancement of agents that operate on web pages. Recent research such as ViperGPT [40] Chameleon [26], RCI Agent [17], VisProg [12], and [28] employ LLMs for planning or action prediction in developing autonomous agents. Benchmark datasets, such as MiniWoB [37], WebShop [51], Macaw-LLM [27], ASH-Prompting [38] Mind2Web [9] and WebArena [57] have also been proposed to measure the ability of LLM-based agents in automating web tasks. These methods mainly involve agents that operate on a text-based Document Object Model (DOM) of HTML scripts. This limits their understanding of screen context, which is crucial for the model’s decision-making and action-taking processes. To address this limitation, Rawles et al. [32] released Android in the Wild, a dataset comprising screens, natural language instructions, and corresponding actions. Following this,  [54] proposed a multimodal model, AutoUI, which is designed to build an agent on the Android in the Wild dataset confined to the Android ecosystem.
@@ 3.1 Task Formulation
Figure 1 illustrates sample tasks and corresponding outputs for three applications within OmniACT: (1) Stocks (MacOS), (2) Apartments.com (web page), and (3) Weather (MacOS). The first column depicts the input image, and the second column shows the natural language task that is to be executed on the current screen. To execute these tasks, a user must accurately perform a series of operations using the mouse and keyboard. For example, to check the rate of change in Google’s stock price over the last month, the mouse has to be moved to the last month and dragged while holding the left-click button to the current month.
@@ 3.2 Dataset Preparation
Table 2: Action type supported by OmniACT and the number of instances for each of the actions in the dataset.
@@ 3.2.2 UI Screen Segmentation
    Desktop Applications: We build a custom annotation interface based on PyQt5111https://pypi.org/project/PyQt5/ to create bounding boxes manually over a screen image using a simple drag-and-click mechanism. This custom interface expedites the process and allows us to get highly accurate gold-label data points for desktop images.
@@
    Websites: For webpages, we write JavaScript code to extract all interactable (click, hover, type, etc.) regions from the HTML source code. We also extract banners, dropdowns, submit, and radio buttons from the screen. We filter the elements to retain only those that are visible and interactable within the screen.
@@ 3.2.3 Functionality Tagging
To map each bounding box to its correct functional description, we leverage Amazon MTurk workers (see details in Appendix), who are given an image with a bounding box and are required to write the correct description or label of the bounding box’s function. For example, given an image of an Amazon webpage with a search bar, the annotator labels it as “find-product-search-bar”. The logical descriptions are used to create tasks in a structured manner without the need to identify individual bounding box coordinates.
@@ 3.2.4 Task Creation
For dataset compilation, college students with basic Python programming skills served as annotators, accessing API references for PyAutoGUI and examples of potential tasks. Each student generated multiple tasks, each accompanied by three alternative natural language reformulations. For instance, “What is 3+2?” might be reformulated as “Calculate the sum of 2 and 3” or “Add two to three”. To avoid train-test leakage, rephrased tasks were consistently placed in the same dataset split. Further details on the annotation process are available in the Appendix.
@@ 3.2.5 Reverse Mapping and Filtering
Table 3: Dataset distribution across splits and platforms.
@@ 4.1 Sequence Score
$SeqScore_{i}=\begin{cases}\beta_{1}+\beta_{2}*(s-1)&\text{if all actions match}\\ 0&\text{otherwise}\end{cases}$
@@ 4.2 Action Score
    Here $L_{2}$ corresponds to the smallest Euclidean distance between the predicted coordinate and bounding box. $L_{2}$ is zero when the predicted coordinate lies within the target bounding box. $\mu$ is the Dirichlet smoothing coefficient which we dynamically set to the length of the diagonal of the bounding box. This ensures that the penalty for points outside the bounding box varies based on the size of the bounding box. For two predicted points with the same $L_{2}$, the metric penalizes more heavily if the box is larger. This is sound with the intuition that the chances of clicking on a larger box are higher and should be penalized more in case of a mistake.
@@
$\alpha_{i}=SeqScore_{i}/\text{ length of sequence $i$}$
@@
Action Score = $\textbf{Action Score = }\\ \frac{\sum_{i}max\left(SeqScore_{i}-\sum_{j}(M_{i}^{j}+K_{i}^{j}+W_{i}^{j}),0\right)}{\sum_{i}SeqScore_{i}}$
We report and discuss these metrics for all baseline models in Section 7.
@@ 5 DetACT: DETecting ACTions from UI
Once all the elements of each category are extracted with their coordinates, we then filter these UI elements by prompting GPT-4 [30]. We ensure that the elements selected are suited only for our task, for which we also provide the task description in our prompts along with the list of elements. Full details of the prompt are provided in the appendix section of the paper. As we observe in Figure 3, given an image from the Expedia application, and a task (“Click on the Black Location icon and enter the destination as Paris.”), the LLM filters out the elements to retain only “Going To”, “Location Icon”, and the Black colored elements from the screen. This is passed as input to the LLM or vision-language model backbone.
@@ 6 Baselines
To evaluate the performance of existing language model-based agents on OmniACT, we conduct experiments with both language-based and multimodal baselines. The DetACT module takes in image and text descriptions of the task and outputs the color, icon, and text-based signals. This is concatenated to the prompt for the LLM prompt-based baselines (see Figure 4). Every prompt starts with a role assignment [55], followed by the detailed API reference of the PyAutoGUI function set, along with a textual description of their function. We then add five in-context examples from the training set that most closely match the task (based on the cosine similarity of the MiniLM [46] embeddings of the reference task and the train examples). We add a list of UI elements filtered by the DetACT module to the prompt. Finally, we provide the rules with the task description. For multimodal baselines, we also pass the image pixels to the vision encoder. We report the results of several baselines:
@@ 7 Results and Analysis
Model Sequence Score (normal-$\uparrow$) Action Score (normal-$\uparrow$)
GPT-4 [30] 36.42 12.77
GPT-4V [50] 39.43 20.76
Table 5: Results of GPT-4 and GPT-4V on a subset of 500 samples.
@@
To address this, we experiment with multimodal language models (Table 4). We observe that the coordinate prediction improves significantly when we provide the entire image as input to the multimodal LLM, as this enables it to fully utilize the screen representation. In addition to open sourced models, we also experiment with the GPT-4-vision-preview API [50] on a subset of 500 data points (due to cost overheads and OpenAI API request limits). Table 5 shows that GPT-4 Vision [50] outperforms GPT-4 significantly on the Action Score along with improving the sequence score, which we attribute to the strong reasoning abilities of GPT-4 coupled with the improved visual understanding capabilities of the GPT-4-vision-preview model [50]. These findings pave the way towards exciting new research directions on building multimodal models for long-horizon planning and code generation.

### v3

# https://huggingface.co/datasets/Writer/omniact

@@ Abstract
footnotetext: * These authors contributed equally. The order is determined by dice rolling.
For decades, human-computer interaction has fundamentally been manual. Even today, almost all productive work done on the computer necessitates human input at every step. Autonomous virtual agents represent an exciting step in automating many of these menial tasks. Virtual agents would empower users with limited technical proficiency to harness the full possibilities of computer systems. They could also enable the efficient streamlining of numerous computer tasks, ranging from calendar management to complex travel bookings, with minimal human intervention. In this paper, we introduce OmniACT, the first-of-a-kind dataset and benchmark for assessing an agent’s capability to generate executable programs to accomplish computer tasks. Our scope extends beyond traditional web automation, covering a diverse range of desktop applications. The dataset consists of fundamental tasks such as “Play the next song", as well as longer horizon tasks such as “Send an email to John Doe mentioning the time and place to meet". Specifically, given a pair of screen image and a visually-grounded natural language task, the goal is to generate a script capable of fully executing the task. We run several strong baseline language model agents on our benchmark. The strongest baseline, GPT-4, performs the best on our benchmark However, its performance level still reaches only 15% of the human proficiency in generating executable scripts capable of completing the task, demonstrating the challenge of our task for conventional web agents. Our benchmark provides a platform to measure and evaluate the progress of language model agents in automating computer tasks and motivates future work towards building multimodal models that bridge large language models and the visual grounding of computer screens.
Keywords:
Generalist Agent Multimodal Machine Learning Dataset and Benchmark Vision Language Understanding UI grounding Human-computer interaction
Figure 1: OmniACT dataset and benchmark for enabling autonomous human-computer interaction agents. The left shows an image paired with a natural language task description as the input, the right shows the resulting action script to be executed on the screen. Examples are presented from Stocks, Apartments, and Weather application.
@@ 1 Introduction
Performing computer tasks based on natural language instructions has been a long-standing goal of artificial intelligence [49]. One concrete objective in the line of research is to develop generalist agents that can assist humans in doing computer tasks [21], such as “Order a pizza from Domino’s" or “Write a message to John." The agent should be able to open the application and perform the task. Executing these actions on a personal computer involves a sequence of interactions with a mouse and keyboard. For example, the simple task of writing an email involves hovering over the application icon, clicking it, clicking the ‘New Email’ button, writing the content of the email, and clicking send. Successfully sending an email requires accurately predicting the correct action at each step and accurately executing it, which is a herculean task even for the best agents today [14].
@@
We evaluate several language model-based agent baselines on this dataset, including LLaMA [47], Vicuna [7], Palmyra-X (43B) [2], InstructPalmyra-30B [45], GPT 3.5, and GPT-4 [32]. We experiment with fine-tuning Vicuna-13B and LLaMA-13B models using QLoRA [10]. We also benchmark multimodal baseline LLaVa-v1.5-7B, LLaVa-v1.5-13B [47], Gemini-Pro [44] and GPT-4-vision-preview [55] for the task. Our findings highlight the necessity for a multimodal model capable of executing these tasks, and our analysis provides insights into promising future work in the space. Our key contributions are outlined as follows:
@@ 2 Related Work
Table 1: Comparison of OmniACT with other related benchmarks.
@@ 2.1 UI Understanding
User interface (UI) understanding has garnered interest from researchers in the machine learning and human-computer interaction communities, evolving with various models focusing on understanding the semantics of mobile and web user interfaces. UIBert [3], PixelBERT [16], ActionBert [15], VUT [25], Screen2Words [48], WidgetCaptioning [24] and Pix2Act [39] are notable models in this area. They propose approaches for learning the user-interface semantics of the mobile screen using the image and view hierarchy. These models have demonstrated effectiveness in tasks like capability prediction, screen segmentation and understanding, and screen caption generation. Lexi [4] and Spotlight [22] propose models that use vision-only inputs to minimize the reliance on metadata such as view hierarchy. Furata et al. [11] demonstrates the use of fine-tuning for multimodal web navigation. The majority of machine learning models trained for UI understanding leverage the Rico dataset [8] and its extensions, which contain 64,462 unique Android screens and metadata. In addition, [4] released the UICaptions dataset, which consists of diverse image-captions pairs across a wide range of applications. PixelHelp [23] also released a corpus to train models that can interpret natural language instructions and map them to mobile UI actions.
@@ 2.2 Autonomous Computer Agents
The advent of large language models (LLMs) has been pivotal in the rapid advancement of agents that operate on web pages. Recent research such as ViperGPT [43] Chameleon [29], RCI Agent [18], VisProg [12], and [31] employ LLMs for planning or action prediction in developing autonomous agents. Benchmark datasets, such as MiniWoB [40], WebShop [56], Macaw-LLM [30], ASH-Prompting [41] Mind2Web [9], WebArena [62], AgentBench [28] and VisualWebArena [20] have also been proposed to measure the ability of LLM-based agents to automate web tasks. These methods mainly involve agents that operate on a text-based Document Object Model (DOM) of HTML scripts. This limits their understanding of screen context, which is crucial for the model’s decision-making and action-taking processes. To address this limitation, [35] released Android in the Wild, a dataset comprising screens, natural language instructions, and corresponding actions. Following this,  [59] proposed a multimodal model, AutoUI, which is designed to build an agent on the Android in the Wild dataset confined to the Android ecosystem. WebAgent [13] utilized Flan-U-PaLM, for grounded code generation, and HTML-T5 and showed improvement on real-world websites.
@@ 3.1 Task Formulation
Figure 1 illustrates sample tasks and corresponding outputs for three applications within OmniACT: (1) Stocks (MacOS), (2) Apartments.com (web page), and (3) Weather (MacOS). The first column depicts the input image, and the second column shows the natural language task that is to be executed on the current screen. To execute these tasks, a user must accurately perform a series of operations using the mouse and keyboard. Eg: to check the rate of change in Google’s stock price over the last month, the mouse has to be moved to the last month and dragged while holding the left-click button to the current month.
@@ 3.2 Dataset Preparation
Table 2: Action types supported by OmniACT and the number of instances for each action in the dataset.
@@ 3.2.2 UI Screen Segmentation
    Desktop Applications: We build a custom annotation interface based on PyQt5111https://pypi.org/project/PyQt5/ to create bounding boxes manually over a screen image using a simple drag-and-click mechanism. This custom interface expedites the process and allows us to get highly accurate gold-label data for desktop images.
@@
    Websites: For webpages, we write JavaScript code to extract all interactable (click, type, etc.) regions from HTML source code. We also extract banners, dropdowns, submit, and radio buttons from the screen. We filter the elements to retain only those that are visible and interactable within the screen.
@@ 3.2.3 Functionality Tagging
To map each bounding box to its correct functional description, we leverage Amazon MTurk workers (see details in Appendix), who are given an image with a bounding box and are required to write the correct description or label of the bounding box’s function. For example, given an image of an Amazon webpage with a search bar, the annotator labels it as “find-product-search-bar". The logical descriptions are used to create tasks in a structured manner without the need to identify individual bounding box coordinates.
@@ 3.2.4 Task Creation
For dataset compilation, college students with basic Python programming skills served as annotators, accessing API references for PyAutoGUI and examples of potential tasks. Each student generated multiple tasks, each accompanied by three alternative natural language reformulations. For instance, “What is 3+2?" might be reformulated as “Calculate the sum of 2 and 3" or “Add two to three". To avoid train-test leakage, rephrased tasks were consistently placed in the same dataset split. Further details on the annotation process are available in the Appendix.
@@ 3.2.5 Reverse Mapping and Filtering
Table 3: Dataset distribution across splits and platforms.
@@ 4.1 Sequence Score
$SeqScore_{i}=\begin{cases}\beta_{1}+\beta_{2}*(s-1)&\text{if all actions match}\\ 0&\text{otherwise}\end{cases}$ (1)
@@ 4.2 Action Score
    Here $L_{2}$ corresponds to the smallest Euclidean distance between the predicted coordinate and bounding box. $L_{2}$ is zero when the predicted coordinate lies within the target bounding box. $\mu$ is the Dirichlet smoothing coefficient which we dynamically set to the inverse of the length of the diagonal of the bounding box. This ensures that the penalty for points outside the bounding box varies based on the size of the bounding box. For two predicted points with the same $L_{2}$, the metric penalizes more heavily if the box is larger. This is sound with the intuition that the chances of clicking on a larger box are higher and should be penalized more in case of a mistake.
@@
$\alpha_{i}=SeqScore_{i}/\text{ length of sequence $i$}$ (5)
@@
Action Score = $\textbf{Action Score = }\frac{\sum_{i}max\left(SeqScore_{i}-\sum_{j}(M_{i}^{j}+K_{i}^{j}+W_{i}^{j}),0\right)}{\sum_{i}SeqScore_{i}}$ (6)
5 DetACT: DETecting ACTions from UI
@@
Once all the elements of each category are extracted with their coordinates, we then filter these UI elements by prompting GPT-4 [32]. We ensure that the elements selected are suited only for our task, for which we also provide the task description in our prompts along with the list of elements. Full details of the prompt are provided in the appendix section of the paper. As we observe in Figure 3, given an image from the Expedia application, and a task (“Click on the Black Location icon and enter the destination as Paris."), the LLM filters out the elements to retain only “Going To", “Location Icon", and the Black colored elements from the screen. This is passed as input to the LLM or VLM backbone.
@@ 6 Baselines
To evaluate the performance of existing language model-based agents, we conduct experiments with both language-based and multimodal baselines. The DetACT module takes in image and text descriptions of the task and outputs the color, icon, and text-based signals. This is concatenated to the prompt for the LLM prompt-based baselines (see Figure 4). Every prompt starts with a role assignment [60], followed by the detailed API reference of the PyAutoGUI function set, along with a textual description of their function. We then add five in-context examples from the training set that most closely match the task (based on the cosine similarity of the MiniLM [50] embeddings of the reference task and the train examples). We add a list of UI elements filtered by the DetACT module to the prompt. Finally, we provide the rules with the task description. For multimodal baselines, we also pass the image pixels to the vision encoder. We choose coordinate-based UI elements in the prompt as recent techniques like the Set-of-Mark (SOM) [54] prompting does not work for desktop settings since it is difficult to obtain interactive elements from the desktop screen images. We report the results of several baselines:
@@ 7 Results and Analysis
To address this, we experiment with multimodal language models (Table 4). We observe that the coordinate prediction improves significantly when we provide the entire image as input to the multimodal LLM, as this enables it to fully utilize the screen representation. In addition to open sourced models, we also experiment with the GPT-4-vision API [55] which shows that GPT-4 Vision [55] outperforms GPT-4 significantly on the Action Score along with improving the sequence score, which we attribute to the strong reasoning abilities of GPT-4 coupled with the improved visual understanding capabilities of the GPT-4-vision model [55]. These findings pave the way towards exciting new research directions on building multimodal models for long-horizon planning and code generation.